# Anomalous dimensions from conformal field theory: generalized $\phi^{2n+1}$ theories

### v2

@@ 2.1 Multiplet recombination
If we divide the left-hand side by those of (13) and (18), then the right-hand side becomes 2, so $\lim_{\epsilon\rightarrow 0}(\epsilon\,\gamma_{\phi}^{-1})$ is finite for positive integer $k$ and $\gamma_{\phi}$ is of order $\epsilon$. It follows that $\alpha$ and $\lambda$ are of order $\epsilon^{1/2}$. Moreover, we derive the explicit expression for the leading anomalous dimension
$\gamma_{\phi}=\frac{\epsilon}{2\left(3+\frac{(-1)^{k}k!\left(3k\right)_{k}}{2^{4k-4}\left(1/2\right)_{k}^{2}}\right)}+O(\epsilon^{3/2})\,,$ (22)
which generalizes the $k=1$ result in Fisher:1978pf . For $k=1$, the anomalous dimension of $\phi$ is negative: $\gamma_{\phi}=-\epsilon/18$, implying that the theory is non-unitary.  [^15] For $k>1$, the generalized free theory itself is non-unitary due to the violation of the unitarity bound. We also obtain
$\alpha^{2}$ $=-\frac{(2k)!\,(3k-1)!\,\epsilon}{2^{5-4k}k!+6(-1)^{k}\frac{\left(1/2\right)_{k}^{2}}{\left(3k\right)_{k}}}+O(\epsilon^{3/2})\,,$ (23)
$\lambda_{\phi\phi\phi}$ $=\frac{(-1)^{k}\alpha}{2^{6k-5}(3/2)_{k-1}^{2}}+O(\epsilon)\,.$ (24)
@@
$4^{2k}(2k-m-1)!\,(-m)_{m}(1-3k-\ell-m)_{2k}\lim_{\epsilon\rightarrow 0}\left[\alpha^{-2}\left(\gamma_{\phi}-\frac{\gamma_{\mathcal{J}}}{2}\right)\right]=\frac{\lambda_{\phi^{2}\phi^{2}\mathcal{J},\text{f}}}{\lambda_{\phi\phi\mathcal{J},\text{f}}}\,,$ (28)
@@
$\gamma_{\mathcal{J}}=2\gamma_{\phi}-\frac{\alpha^{2}}{2^{4k-3}(2k-m-1)!(-m)_{m}(1-3k-\ell-m)_{2k}}+O(\epsilon^{3/2})\,,$ (29)
where $\gamma_{\phi}$ and $\alpha^{2}$ are given in (22) and (23). This is the $k\geqslant 1$ generalization of the $k=1$ result in Giombi:2016hkj . The stress tensor is the spin-2 current on the highest trajectory $m=k-1$, and its anomalous dimension vanishes to order $\epsilon$, according to (29).
The spin-0 operator on the lowest trajectory $\mathcal{J}_{\ell=0}^{(m=0)}=\phi^{2}$ is not a primary, so the matching condition (25) does not apply in this case. Nevertheless, we can set $\ell=m=0$ in (29). It turns out that we obtain the shadow of $\phi$:
$\Delta_{\smash{\phi}}+\Delta_{\mathcal{J}}\big{|}_{\ell=m=0}=d+O(\epsilon^{3/2})\,,$ (30)
@@ 2.2 Analytic bootstrap
In this subsection, we verify the consistency between the results above and crossing symmetry. In the lightcone limit of the crossing equation for the four-point function $\langle\phi\phi\phi\phi\rangle$, the anomalous dimensions of the bilinear operators are read off from the terms with enhanced singularity in the crossed channel. [^17] The meaning of the enhanced singularity is as follows. An expression with an enhanced singularity is power-law divergent after repeated action of the conformal Casimir operator (see e.g. Alday:2015eya ). The results are consistent with those derived from the multiplet recombination.
@@
We expand the left-hand side of (33) in the anomalous dimension
@@
where we have suppressed the indices associated with $\mathcal{J}_{\ell}^{(0)}$ for simplicity. Consider the right-hand side of the crossing equation (33) at leading order in the lightcone expansion. The identity operator contributes to the enhanced singularity at order $\epsilon^{0}$. The corresponding term is $(\frac{1-\bar{z}}{\bar{z}})^{-\Delta_{\smash{\phi_{\text{f}}}}}$, with $\Delta_{\smash{\phi_{\text{f}}}}=2k$. The free squared OPE coefficients of $\mathcal{J}^{(m=0)}_{\ell}$ can be derived using (39):
@@
where $\mathcal{O}$ is a scalar operator. We focus only on the terms with the enhanced singularity at order $\epsilon^{1}$:
@@
where the Casimir-regular terms do not contribute to the enhanced singularity. Using (24), (37), (39) and (43), we obtain the anomalous dimensions $\tilde{\gamma}_{\mathcal{J}}$. The result agrees with (29), confirming the consistency of the multiplet-recombination results (24) and (29) with crossing symmetry.
@@
$\lim_{\epsilon\rightarrow 0}\left[P_{\phi}G_{\phi}(1-\bar{z},1-z)\Big{|}_{\log z\text{ part}}\,\right]=P_{\phi^{2}_{\text{f}}}\,G_{\phi^{2}_{\text{f}}}(1-\bar{z},1-z)\Big{|}_{\log z\text{ part}}\,,$ (46)
which is compatible with the multiplet recombination.
@@ 3 Generalized Potts model
We can extend the $\phi^{3}$ discussion to the case of an $N$-component field $\phi_{a}$, where $a=1,2,\ldots,N$. The action that describes the generalized ($N+1$)-state Potts model is
$S=\int\mathrm{d}^{d}x\left(\sum_{a=1}^{N}\phi_{a}\Box^{k}\phi_{a}+g\mu^{\epsilon/2}\sum_{a,b,c=1}^{N}T_{abc}\,\phi_{a}\phi_{b}\phi_{c}\right)\,,$ (47)
where $T_{abc}$ is an $S_{N+1}$ invariant tensor [^18] The rank-3 invariant tensor of $S_{N+1}$ is unique up to a scalar multiple wallace1978spin .
@@
In this section, we obtain the anomalous dimension of $\phi_{a}$ and bilinear operators $\mathcal{J}_{\ell}^{(m)}$, using the multiplet recombination. For positive integers $N>2$, there are four types of bilinear operators, due to the tensor product decomposition murnaghan1938analysis
@@
We refer to these four representations as the trivial, standard, symmetric traceless and antisymmetric representations (see appendix C for an explanation of these names). The general results for the anomalous dimensions depend on $N$, and some particularly interesting cases are the percolation theory ($N\rightarrow 0$) and the spanning forest ($N\rightarrow-1$). As in subsection 2.2, we verify the recombination results using the analytic bootstrap.
@@ 3.1 Multiplet recombination
The multi-field version of (6) takes the form
@@
The left-hand side of (51) is nearly identical to that of (8). The difference is that here we have a tensor structure $\delta_{a_{1}a_{2}}$, which is the only rank-2 $S_{N+1}$-invariant tensor (up to a scalar multiple). [^19] The uniqueness can be seen from the tensor product decomposition (49), where the trivial representation occurs only once on the right-hand side wallace1978spin . The correlator on the right-hand side is again given by Wick’s theorem. We obtain
@@
which results in
$\lim_{\epsilon\rightarrow 0}\Bigg{\{}\alpha^{-3}$ $\sum^{k}_{\begin{subarray}{c}i_{1},j_{1}=0\\ i_{1}+j_{1}\geqslant k\end{subarray}}\;\;\sum^{k}_{\begin{subarray}{c}i_{2},j_{2}=0\\ i_{2}+j_{2}\geqslant k\end{subarray}}\;\;\sum^{k}_{\begin{subarray}{c}i_{3},j_{3}=0\\ i_{3}+j_{3}\geqslant k\end{subarray}}\Bigg{[}B_{i_{1},j_{1}}C_{i_{1},j_{1}}^{\Delta_{\smash{\phi}},\Delta_{\smash{\phi}}}B_{i_{2},j_{2}}C_{i_{2},j_{2}}^{\Delta_{\smash{\phi}}+2i_{1},\Delta_{\smash{\phi}}+2(k-i_{1}-j_{1})}$
×$\!\!\times B_{i_{3},j_{3}}C_{i_{3},j_{3}}^{\Delta_{\smash{\phi}}+2(j_{2}+k-i_{1}-j_{1}),\Delta_{\smash{\phi}}+2(j_{1}+k-i_{2}-j_{2})}x_{12}^{2(i_{3}+j_{3}-i_{1}-i_{2}-k)}$
×$\!\!\times x_{13}^{2(i_{2}+j_{2}-j_{1}-j_{3}-k)}x_{23}^{2(i_{1}+j_{1}-j_{2}-i_{3}-k)}\Bigg{]}\frac{\lambda_{\phi\phi\phi}T_{a_{1}a_{2}a_{3}}}{x_{12}^{\Delta_{\smash{\phi}}}x_{13}^{\Delta_{\smash{\phi}}}x_{23}^{\Delta_{\smash{\phi}}}}\Bigg{\}}=\frac{8(N-2)(N+1)^{2}\,T_{a_{1}a_{2}a_{3}}}{x_{12}^{\Delta_{\smash{K_{\text{f}}}}}x_{13}^{\Delta_{\smash{K_{\text{f}}}}}x_{23}^{\Delta_{\smash{K_{\text{f}}}}}}\,.$ (59)
In the $\epsilon$ expansion, this leads to the constraint
@@
$\gamma_{\phi}=\frac{\epsilon}{2\left(3+\frac{N-2}{N-1}\frac{(-1)^{k}k!(3k)_{k}}{2^{4k-4}(1/2)_{k}^{2}}\right)}+O(\epsilon^{3/2})\,.$ (61)
This expression is very similar to (22), but here an $N$-dependent factor appears in the second term in the denominator. [^20] Interestingly, the anomalous dimension of $\phi$ is $k$-independent to order $\epsilon$ at $N=2$: $\gamma_{\phi}=\epsilon/6+O(\epsilon^{3/2})$. For $N=1$, we have $\gamma_{\phi}=O(\epsilon^{3/2})$ due to $T_{111}=0$. See Fang:2022ufx ; Fang:2022tav and references therein for the Fortuin-Kasteleyn representation of the Ising model above the standard upper critical dimension, i.e., $d>4$. At $k=1$, the result agrees with that from the RG analysis Priest:1976zz ; Amit:1976pz . We also obtain $\alpha$ and the OPE coefficient of $\phi$:
$\alpha^{2}$ $=-\frac{(2k)!\,(3k-1)!\,\epsilon}{(N+1)^{2}\left((N-2)2^{5-4k}k!+6(N-1)(-1)^{k}\frac{(1/2)_{k}^{2}}{(3k)_{k}}\right)}+O(\epsilon^{3/2})\,,$ (62)
$\lambda_{\phi\phi\phi}$ $=\frac{(-1)^{k}\alpha}{2^{6k-5}(3/2)_{k-1}^{2}}+O(\epsilon^{3/2})\,.$ (63)
As mentioned in the introduction, the $N\rightarrow\infty$ limit yields the decoupled Yang-Lee theory. One can check that
@@
$\gamma_{\phi}\big{|}_{N\rightarrow 0}=\frac{\epsilon}{2\left(3+\frac{(-1)^{k}k!(3k)_{k}}{2^{4k-5}(1/2)_{k}^{2}}\right)}+O(\epsilon^{3/2})\,,\qquad\gamma_{\phi}\big{|}_{N\rightarrow-1}=\frac{\epsilon}{2\left(3+\frac{3(-1)^{k}k!(3k)_{k}}{2^{4k-3}(1/2)_{k}^{2}}\right)}+O(\epsilon^{3/2})\,.$ (65)
At $k=1$, they lead to the critical exponents of percolation theory $\eta_{\text{percolation}}=-\epsilon/21+O(\epsilon^{3/2})$ and spanning forest $\eta_{\text{spanning forest}}=-\epsilon/15+O(\epsilon^{3/2})$.
@@
where $\lambda_{KK\mathcal{J},\text{f}}$ denotes the OPE coefficient of $\mathcal{J}_{ab,\text{f}}\in K_{a,\text{f}}\times K_{b,\text{f}}$. The tensor structures $\mathbf{t}$ on both hand sides match, because the operators on both hand sides of (66) live in the same $S_{N+1}$ representations. The constraint on the anomalous dimensions can be written collectively as
$2^{4k}(2k-m-1)!\,(-m)_{m}(1-3k-\ell-m)_{2k}\lim_{\epsilon\rightarrow 0}\left[\alpha^{-2}\left(\gamma_{\phi}-\frac{\gamma_{\mathcal{J}}}{2}\right)\right]=\frac{\lambda_{KK\mathcal{J},\text{f}}}{\lambda_{\phi\phi\mathcal{J},\text{f}}}\,,$ (68)
which is of the same form as (28). The difference is that here the ratio of free OPE coefficients depend on $N$. The result is
$\gamma_{\mathcal{J}}=2\gamma_{\phi}-\frac{R\,\alpha^{2}}{2^{4k-1}(2k-m-1)!(-m)_{m}(1-3k-\ell-m)_{2k}}+O(\epsilon^{3/2})\,,$ (69)
@@
For the spin-0 trivial and symmetric traceless operators in the canonical case, the results agree with theumann1984crossover . On the highest trajectory $m=k-1$, the spin-2 current in the trivial representation $(N+1)$ is the stress tensor. As expected, the anomalous dimension of the stress tensor is $O(\epsilon^{3/2})$ according to (69). At $k=1$, the set of leading anomalous dimensions $\{\gamma_{\phi},\gamma_{\mathcal{J}}\}$ contains negative elements when $N>7/3$, which indicates that the canonical Potts model is non-unitary near the upper critical dimension. Moreover, there exists a shadow relation in the standard representation:
$\Delta_{\smash{\phi}}+\Delta_{\mathcal{J}_{\text{standard}}}\big{|}_{\ell=m=0}=d+O(\epsilon^{3/2})\,,$ (71)
which is true for any positive integer $k$. In the $N\rightarrow\infty$ limit, this equation is related to (30) in the generalized Yang-Lee theory. If we take the limit $N\rightarrow 0$, there are degeneracies in the scaling dimensions, leading to logarithmic structures (see appendix G). We also discuss some special limits of the generalized O($N$)-symmetric $\phi^{2n}$ theory in appendix H. In the limit $N\rightarrow-1$, the ratio of OPE coefficients (70) vanishes, but the formal product $R\,\alpha^{2}$ in (69) remains finite because $\alpha^{2}$ diverges as $1/(N+1)^{2}$.
Below, we examine the Chew-Frautschi plots to gain further insights into the theory. According to the procedure in Caron-Huot:2022eqs , the analytic continuation of the trajectories is given by
@@
=\;= $\left(k+2m+\ell+2\gamma_{1}-\frac{R\,\alpha^{2}}{2^{4k-1}(2k-m-1)!(1-3k-m-\ell)_{2k}(-m)_{m}}-\frac{\epsilon}{2}+O(\epsilon^{3/2})\right)^{2}\,.$ (72)
@@
Figure 1: Chew-Frautschi plot for $\mathcal{J}^{(m=0)}_{\ell,\,\text{trivial}}$ at $N=3$. The dashed lines indicate the $\epsilon\rightarrow 0$ limit of the Regge trajectory. We have set $k=1$ and $\epsilon=10^{-3}$.
@@
Figure 3: Regge trajectories associated with $\mathcal{J}^{(m=0)}_{\ell,\,\text{trivial}}$. The trajectory in the interacting theory approaches the free theory trajectory as $N\rightarrow 1$. We have used $k=1$ and $\epsilon=10^{-3}$.
@@
$\Delta^{*}_{\pm}-\frac{d}{2}=\pm\left(1+\frac{3R^{\prime}-11N+23}{6(3N-7)}\,\epsilon+O(\epsilon^{3/2})\right)\,,$ (73)

### v3

@@ 2.1 Multiplet recombination
If we divide the left-hand side by those of (13) and (16), then the right-hand side becomes 2. We have
$\frac{(-1)^{k}2^{4k-4}\left(1/2\right)_{k}^{2}}{k!\left(3k\right)_{k}}\left[\,\lim_{\epsilon\rightarrow 0}\left(\epsilon\gamma_{\phi}^{-1}\right)-6\right]=2\,.$ (19)
A special case occurs if the coefficient outside the square brackets is $-1/3$. However, this is impossible for positive integer $k$. The reason is as follows. The absolute value of the coefficient increases with $k$. This absolute value is smaller than $1/3$ for $k=4$, and slightly larger than $1/3$ for $k=5$. This means that the coefficient cannot be $-1/3$. We conclude that $\lim_{\epsilon\rightarrow 0}(\epsilon\,\gamma_{\phi}^{-1})$ is finite for positive integer $k$ and $\gamma_{\phi}$ is of order $\epsilon$. It follows that $\alpha$ and $\lambda$ are of order $\epsilon^{1/2}$. Moreover, we derive the explicit expression for the leading anomalous dimension
$\gamma_{\phi}=\frac{\epsilon}{2\left(3+\frac{(-1)^{k}k!\left(3k\right)_{k}}{2^{4k-4}\left(1/2\right)_{k}^{2}}\right)}+\ldots\,,$ (20)
which generalizes the $k=1$ result in Fisher:1978pf . Here and below, we use the ellipsis to denote the higher-order terms in $\epsilon$.  [^17] In the canonical case $k=1$, the higher-order terms in $\gamma_{\phi}$ are known to be associated with integer powers of $\epsilon$ deAlcantaraBonfim:1981sy . It is natural to expect that only integer powers of $\epsilon$ appear in the generic $k>1$ counterparts as well. To verify this, it would be interesting to carry out the explicit computation for the $k>1$ theories in the more standard diagrammatic approach. For $k=1$, the anomalous dimension of $\phi$ is $\gamma_{\phi}=-\epsilon/18+O(\epsilon^{2})$, which is negative and implies that the canonical Yang-Lee theory is nonunitary. [^18] For $k>1$, the generalized free theory itself is nonunitary due to the violation of the unitarity bound. The anomalous dimension $\gamma_{\phi}$ is particularly large for $k=5$, as the denominator in (20) is small in this case. In addition, we also obtain
$\alpha^{2}$ $=-\frac{(2k)!\,(3k-1)!\,\epsilon}{2^{5-4k}k!+6(-1)^{k}\frac{\left(1/2\right)_{k}^{2}}{\left(3k\right)_{k}}}+\ldots\,,$ (21)
$\lambda_{\phi\phi\phi}$ $=\frac{(-1)^{k}\alpha}{2^{6k-5}(3/2)_{k-1}^{2}}+\ldots$ (22)
@@
$(-1)^{m}4^{2k}(2k-m-1)!m!(k+\ell+m)_{2k}\lim_{\epsilon\rightarrow 0}\left[\alpha^{-2}\left(\gamma_{\phi}-\frac{\gamma_{\mathcal{J}}}{2}\right)\right]=\frac{\lambda_{\phi^{2}\phi^{2}\mathcal{J},\text{f}}}{\lambda_{\phi\phi\mathcal{J},\text{f}}}\,,$ (26)
@@
$\gamma_{\mathcal{J}^{(m)}_{\ell}}=2\gamma_{\phi}-\frac{(-1)^{m}\alpha^{2}}{2^{4k-3}m!(2k-m-1)!(k+\ell+m)_{2k}}+\ldots\,,$ (27)
where $\gamma_{\phi}$ and $\alpha^{2}$ are given in (20) and (21). This is the $k\geqslant 1$ generalization of the $k=1$ result in Giombi:2016hkj . The stress tensor is the spin-2 current on the highest trajectory $m=k-1$, and its anomalous dimension vanishes to order $\epsilon$, as expected.
The spin-0 operator on the lowest trajectory $\mathcal{J}_{\ell=0}^{(m=0)}=\phi^{2}$ is not a primary, so the matching condition (23) does not apply in this case. Nevertheless, we can set $\ell=m=0$ in (27). It turns out that we obtain the shadow of $\phi$,
$\Delta_{\smash{\phi}}+\Delta_{\mathcal{J}}\big{|}_{\ell=m=0}=d+0\epsilon+\ldots\,,$ (28)
@@ 2.2 Analytic bootstrap
In this subsection, we verify the consistency between the results above and crossing symmetry. In the lightcone limit of the crossing equation for the four-point function $\langle\phi\phi\phi\phi\rangle$, the anomalous dimensions of the bilinear operators are read off from the terms with an enhanced singularity in the crossed channel. [^20] The meaning of the enhanced singularity is as follows. An expression with an enhanced singularity is power-law divergent after repeated action of the conformal Casimir operator (see, e.g., Alday:2015eya ). The results are consistent with those derived from the multiplet recombination.
@@
Let us expand the left-hand side of (31) in the anomalous dimension
@@
where we have suppressed the indices associated with $\mathcal{J}_{\ell}^{(0)}$ for simplicity. Consider the right-hand side of the crossing equation (31) at leading order in the lightcone expansion. The identity operator contributes to the enhanced singularity at order $\epsilon^{0}$. The corresponding term is $(\frac{1-\bar{z}}{\bar{z}})^{-\Delta_{\smash{\phi_{\text{f}}}}}$, with $\Delta_{\smash{\phi_{\text{f}}}}=2k$. The free squared OPE coefficients of $\mathcal{J}^{(0)}_{\ell}$ can be derived using (37):
@@
where $\mathcal{O}$ is a scalar operator. We have
@@
where the Casimir-regular terms do not contribute to the enhanced singularity. Using the OPE coefficient (22) derived from the multiplet recombination, we obtain the anomalous dimension
$\tilde{\gamma}_{\mathcal{J}_{\ell}^{(0)}}=-\frac{\alpha^{2}}{2^{4k-3}(2k-1)!(k+\ell)_{2k}}+\ldots\,,$ (44)
which agrees with (27) at $m=0$, confirming the consistency of the multiplet-recombination results (22) and (27) with crossing symmetry. Note that this agreement holds for arbitrary $\alpha$ [^21] In the generalized $\phi^{2n}$ theories, we used the analytic bootstrap to verify the multiplet-recombination results as well Guo:2023qtt . The results from the two approaches also agree for arbitrary $\alpha$.
@@
$\lim_{\epsilon\rightarrow 0}\left[P_{\phi}G_{\phi}(1-\bar{z},1-z)\right]=P_{\phi^{2}_{\text{f}}}\,G_{\phi^{2}_{\text{f}}}(1-\bar{z},1-z)\,.$ (45)
This is a manifestation of the multiplet recombination in the conformal block expansion.
@@ 3 Generalized Potts model
We can extend the $\phi^{3}$ discussion to the case of an $N$-component field $\phi_{a}$, where $a=1,2,\ldots,N$. The action for the generalized ($N+1$)-state Potts model is
$S\propto\int\mathrm{d}^{d}x\left(\sum_{a=1}^{N}\phi_{a}\Box^{k}\phi_{a}+g\mu^{\epsilon/2}\sum_{a,b,c=1}^{N}T_{abc}\,\phi_{a}\phi_{b}\phi_{c}\right)\,,$ (46)
where $T_{abc}$ is an $S_{N+1}$-invariant tensor [^22] The rank-3 invariant tensor of $S_{N+1}$ is unique up to a scalar multiple wallace1978spin .
@@
In this section, we obtain the anomalous dimension of $\phi_{a}$ and bilinear operators $\mathcal{J}_{\ell}^{(m)}$ using the multiplet recombination. For positive integers $N>2$, there are four types of bilinear operators, due to the tensor product decomposition murnaghan1938analysis
@@
We refer to these four representations as the trivial, standard, symmetric traceless and antisymmetric representations (see appendix C for an explanation of these names). The general results for the anomalous dimensions depend on $N$, and some particularly interesting cases are the percolation theory ($N\rightarrow 0$) and the spanning forest ($N\rightarrow-1$). As in section 2.2, we verify the recombination results using the analytic bootstrap in section 3.2.
@@ 3.1 Multiplet recombination
The multifield version of (6) takes the form
@@
The left-hand side of (50) is nearly identical to that of (8). The difference is that here we have a tensor structure $\delta_{a_{1}a_{2}}$, which is the only rank-2 $S_{N+1}$-invariant tensor (up to a scalar multiple). [^23] The uniqueness can be seen from the tensor product decomposition (48), where the trivial representation occurs only once on the right-hand side wallace1978spin . The correlator on the right-hand side is again given by Wick contractions. We obtain
@@
Again, the action of the Laplacians is given in (A). In the $\epsilon$ expansion, this leads to the constraint
@@
$\gamma_{\phi}=\frac{\epsilon}{2\left(3+\frac{N-2}{N-1}\frac{(-1)^{k}k!(3k)_{k}}{2^{4k-4}(1/2)_{k}^{2}}\right)}+\ldots$ (59)
At $k=1$, the result agrees with that from the renormalization group analysis Priest:1976zz ; Amit:1976pz . We also obtain $\alpha$ and the OPE coefficient of $\phi$:
$\alpha^{2}$ $=-\frac{(2k)!\,(3k-1)!\,\epsilon}{(N+1)^{2}\left((N-2)2^{5-4k}k!+6(N-1)(-1)^{k}\frac{(1/2)_{k}^{2}}{(3k)_{k}}\right)}+\ldots\,,$ (60)
$\lambda_{\phi\phi\phi}$ $=\frac{(-1)^{k}\alpha}{2^{6k-5}(3/2)_{k-1}^{2}}+\ldots\,.$ (61)
The expression (59) is very similar to (20), but here an $N$-dependent factor appears in the second term in the denominator. [^24] Interestingly, the anomalous dimension of $\phi$ is $k$-independent to order $\epsilon$ at $N=2$: $\gamma_{\phi}=\epsilon/6+\ldots$ For $N=1$, we have $\gamma_{\phi}=0\epsilon+\ldots$ due to $T_{111}=0$. See Fang:2022ufx ; Fang:2022tav and references therein for the Fortuin-Kasteleyn representation of the Ising model above the standard upper critical dimension, i.e., $d>4$. As mentioned in the Introduction, the $N\rightarrow\infty$ limit yields the decoupled Yang-Lee theory. One can check that
@@
$\gamma_{\phi}\big{|}_{N\rightarrow 0}=\frac{\epsilon}{2\left(3+\frac{(-1)^{k}k!(3k)_{k}}{2^{4k-5}(1/2)_{k}^{2}}\right)}+\ldots\,,\qquad\gamma_{\phi}\big{|}_{N\rightarrow-1}=\frac{\epsilon}{2\left(3+\frac{3(-1)^{k}k!(3k)_{k}}{2^{4k-3}(1/2)_{k}^{2}}\right)}+\ldots\,.$ (63)
At $k=1$, they lead to the critical exponents of percolation theory $\eta_{\text{percolation}}=-\epsilon/21+O(\epsilon^{2})$, and spanning forest $\eta_{\text{spanning forest}}=-\epsilon/15+O(\epsilon^{2})$.
Note that (59) is divergent for $N=N_{\text{c}}$, where the critical value is
$N_{\text{c}}=1+\frac{1}{1+\frac{3(-1)^{k}2^{4k-4}(1/2)_{k}^{2}}{k!(3k)_{k}}}\,.$ (64)
For $k=1$, we have $N_{\text{c}}=7/3$. In fact, the constraints (53), (56), and (58) have no interacting solution at $N=N_{\text{c}}$ if $\gamma_{\phi}$ is of order $\epsilon^{1}$ or higher. The Potts analog of (19) is
$\lim_{\epsilon\rightarrow 0}\left(\epsilon\gamma_{\phi}^{-1}\right)=\left(6+\frac{(-1)^{k}2^{3+2k}k!(k+1/2)_{k}}{(k)_{2k}}\right)\frac{N-N_{\text{c}}}{N-1}\,,$ (65)
so the way out is to assume that $\gamma_{\phi}$ is of lower order than $\epsilon^{1}$.  [^25] Below (19), we argue that $\gamma_{\phi}$ should be of first order in $\epsilon$. The Yang-Lee result does not apply to the discussion here due to the important finite-$N$ effects, which are not captured by the large $N$ limit associated with the Yang-Lee case. Then we can omit the $\epsilon$ term in (58). In this way, (53), (56), and (58) are consistent with an interacting solution $\alpha\neq 0$. However, as one of the three constraints becomes redundant, there remains a free parameter in $(\gamma_{\phi},\alpha^{2},\lambda_{\phi\phi\phi})$. According to (53), $\alpha^{2}$ is also of lower order than $\epsilon^{1}$ at $N=N_{\text{c}}$. If $\alpha$ takes the form $(-\epsilon)^{\theta}$ with $\theta<\frac{1}{2}$, then this coefficient can have both real and imaginary parts, which is different from a purely imaginary number $(-\epsilon)^{1/2}$ in the canonical Yang-Lee case.  [^26] We assume that the implicit coefficient of $(-\epsilon)^{\theta}$ is a real number. A complex $\alpha$ is associated with a complex CFT, rather than a real or an imaginary CFT.  [^27] As a basic example of imaginary CFTs, the canonical Yang-Lee CFT can be formulated as a real CFT using the rotation $\phi\rightarrow i\phi$. For the canonical case $k=1$, we can make use of the explicit results from the diagrammatic calculation. The two-loop $\beta$ function from deAlcantaraBonfim:1981sy implies that the leading fixed-point coupling constant $g$ takes the form $(-\epsilon)^{1/4}$, so $\alpha\sim g$ is indeed complex. In addition, the complex CFTs can arise for other values of $N$ at lower $d$. We refer to Wiese:2023vgq and references therein for more details on the canonical case $k=1$.
@@
where $\lambda_{KK\mathcal{J},\text{f}}$ denotes the OPE coefficient of $\mathcal{J}_{ab,\text{f}}\in K_{a,\text{f}}\times K_{b,\text{f}}$. The tensor structures $\mathbf{t}$ on both sides match, because the operators on both sides of (66) live in the same $S_{N+1}$ representations. The constraint on the anomalous dimensions can be written collectively as
$(-1)^{m}4^{2k}m!(2k-m-1)!(k+\ell+m)_{2k}\lim_{\epsilon\rightarrow 0}\left[\alpha^{-2}\left(\gamma_{\phi}-\frac{\gamma_{\mathcal{J}}}{2}\right)\right]=\frac{\lambda_{KK\mathcal{J},\text{f}}}{\lambda_{\phi\phi\mathcal{J},\text{f}}}\,,$ (68)
which is of the same form as (26). The difference is that here the ratio of free OPE coefficients depends on $N$. The result is
$\gamma_{\mathcal{J}^{(m)}_{\ell}}=2\gamma_{\phi}-\frac{(-1)^{m}R\,\alpha^{2}}{2^{4k-1}m!(2k-m-1)!(k+\ell+m)_{2k}}+\ldots\,,$ (69)
@@
For the trivial and standard representations, the $N\rightarrow\infty$ limit of (69) leads to the anomalous dimension (27) in the Yang-Lee theory. For the spin-0 trivial and symmetric traceless operators in the canonical case, the results agree with theumann1984crossover . On the highest trajectory $m=k-1$, the spin-2 current in the trivial representation $(N+1)$ is the stress tensor. As expected, the anomalous dimension of the stress tensor is $0\epsilon+\ldots$ according to (69). At $k=1$, the set of leading anomalous dimensions $\{\gamma_{\phi},\gamma_{\mathcal{J}}\}$ contains negative elements when $N>7/3$, which indicates that the canonical Potts model is nonunitary near the upper critical dimension. Moreover, there exists a shadow relation in the standard representation,
$\Delta_{\smash{\phi}}+\Delta_{\mathcal{J}_{\text{standard}}}\big{|}_{\ell=m=0}=d+0\epsilon+\ldots\,,$ (71)
which applies to any positive integer $k$. In the $N\rightarrow\infty$ limit, this equation is related to (28) in the generalized Yang-Lee theory. If we take the limit $N\rightarrow 0$, there are degeneracies in the scaling dimensions, leading to logarithmic structures (see appendix G). [^28] We also discuss some special limits of the generalized O($N$)-symmetric $\phi^{2n}$ theory in appendix H. In the limit $N\rightarrow-1$, the ratio of OPE coefficients (70) vanishes, but the formal product $R\,\alpha^{2}$ in (69) remains finite because $\alpha^{2}$ diverges as $1/(N+1)^{2}$.
Below, we examine the Chew-Frautschi plots to gain further insights into the Regge trajectories of the (generalized) Potts model. According to the procedure in Caron-Huot:2022eqs , the analytic continuation of the trajectories is given by
@@
=\;= $\left(k+2m+\ell+2\gamma_{1}-\frac{(-1)^{m}R\,\alpha^{2}}{2^{4k-1}m!(2k-m-1)!(k+m+\ell)_{2k}}-\frac{\epsilon}{2}+\ldots\right)^{2}\,.$ (72)
@@
Figure 2: Regge trajectories associated with $\mathcal{J}^{(m=0)}_{\ell,\,\text{trivial}}$. The trajectory in the interacting theory approaches the free-theory trajectory as $N\rightarrow 1$. We used $k=1$ and $\epsilon=10^{-3}$.
@@
$\Delta^{*}_{\pm}-\frac{d}{2}=\pm\left(1+\frac{3R^{\prime}-11N+23}{6(3N-7)}\,\epsilon+\ldots\right)\,,$ (73)